OpenStudy (anonymous):

what is 4 minus v over 5 equals 0

OpenStudy (anonymous):

the answer to this question is "I have no Idea!"

OpenStudy (anonymous):

Someone award me please!!!!

OpenStudy (anonymous):

4- (v/5)=0. Move the equation around to form -4=(-v/5) . Then simplify to v equals 20

OpenStudy (anonymous):

But you don't know that if it's (4-v)/5 = 0 Or if it is 4-(v/5) = 0

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.